OpenStudy (anonymous):

The area of a rectangular piece of land is 280 square meters. If the length of the land was 5 meters less and the width was 1 meter more, the shape of the land would be a square. Write an equation to find the width (x) of the land. Show the steps of your work. What is the width of the land in meters? Show the steps of your work.

OpenStudy (anonymous):

Here is what I have for part a l * w = 280 w = 280/l I don't know what the rest is

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Ummm... honestly on these I guess and check, but it doesn't look like you can do that here, give me just a min

OpenStudy (sleepyjess):

l - 5 = w + 1

OpenStudy (anonymous):

yeah thats another thing that i did but i put it as part b, I don't know why

OpenStudy (anonymous):

@aaronq @SolomonZelman

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@perl

OpenStudy (sleepyjess):

I feel like once someone just gets us started, it won't be a problem